Tentative Schedule:

Tuesday  Thursday

Jan 12th 1

1.1, 1.2 Review 14th 2 1.3, 1.4 Review

19th 3 2.1, 2.2 Linear Equations

21st 4

26th  5 28th  6

Feb 2nd  7 4th  8

9th  9 11th  10

16th  11 18th  12

23rd  13 25th  14

Mar 2nd  15 4th  16

9th  17 11th  18

16th  19 18th  20

23rd  21 25th  22

30th  23 Apr 1st  24

6th  25 8th  26

13th  27 15th  28

Tuesday  Thursday
20th  29 22nd  30

27th Final Exam 29th